Find all solutions to the following system of linear equations.

x1 − x2 − x3 = 2

2x1 + x2 + 2x3 = 4

x1 − 4x2 − 5x3 = 2


1
Expert's answer
2021-02-02T14:54:19-0500

X1-X2-X3=2

2X1+X2+2X3=4

X1-4X2-5X3=2

Making X1 the subject of the formula

X1=2+X2+X3

2(2+X2+X3) +X2+2X3=4

4+2X2+2 X3+ X2+2X3=4

Collecting the like terms

3 X2+4 X3=4-4

3X2=-4X3

X3=-4/3X3...(i)

2+X2+X3-4X2-5X3=2

Collecting the like terms

X3-5X3-4X2+ X2=2-2

-4 X3-3 X2=0

3X2=-4X3

X2=-4/3X3...(ii)

∴ -4/3X3=-4/3X3

X3=0 and X2=0

X1=2+0+0

X1=2

∴ X1=2, X2=0, and X3=0